OpenStudy (anonymous):

Evaluate the expression below when x = -1 2x^2 - 5 over x+9 A) -7/8 B) -3/8 C) -1/8 D) -1

OpenStudy (anonymous):

I tried, can you please tell me the answer.. I'm not good at math

OpenStudy (phi):

First, what exactly is the question? is it (2x^2 - 5)/ (x+9) ? i.e. is it \[ \frac{ 2x^2 - 5}{ x+9} \]

OpenStudy (anonymous):

The second one @phi I just never understood how to write

OpenStudy (phi):

if you don't use the math equation editor (button on the lower left of where you type in) then you should put in parens (like I did up above the "nice" version) otherwise it means what Akash posted (which is different)

OpenStudy (phi):

to answer your question, look for the x's in the expression. *erase them* and put in -1 instead. can you do that ?

OpenStudy (anonymous):

study that

OpenStudy (anonymous):

@phi Can you tell me the answer. I'm sick and have a bad headache but I got to get this done :( Sorry for being so blunt

OpenStudy (anonymous):

@iGreen can you help me

OpenStudy (phi):

look for the x's in the expression. *erase them* and put in -1 instead. can you do that ?

OpenStudy (igreen):

\(\dfrac{ 2x^2 - 5}{ x+9}\) Input -1 for all values of x: \(\dfrac{ 2(-1)^2 - 5}{ -1+9}\) Simplify Exponent: \(\dfrac{ 2(1) - 5}{ -1+9}\) Multiply: \(\dfrac{ 2 - 5}{ -1+9}\) Add & Subtract: \(-\dfrac{3}{8}\)

OpenStudy (anonymous):

Your a life saver c:

OpenStudy (igreen):

Lol, np.